Sally bought three chocolate bars and a pack of gum ad paid 1.75. Jake bought two chocolate bars and four packs of gum and paid 2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by mszlmb(115)

Sally bought three chocolate bars and a pack of gum ad paid 1.75. Jake bought
two chocolate bars and four packs of gum and paid 2.00. Find the cost of a
chocolate bar and the cost of a pack of gum. 3C+G=1.75 2C+4G=2 -3c on the first equation G=1.75-3C Replace G in the second equation with its equivalent, 1.75-3C 2C+4(1.75-3C)=2 SIMPLIFY by DISTRIBUTION 2C+7.00-12C=2 -7.00 BOTH SIDES -10C=-5 divide by -10, both sides C=.50 A chocolate bar costs 50 cents, or $.50 Now plug C in to the second equation 2(.5)+4G=2 Simplify 1+4G=2 -1 both sides 4G=1 divide by 4 G=.25 A pack of gum costs 25 cents, or $.25 The answers: A chocolate bar costs $0.50 A pack of gum costs $0.25